# Greedy Algorithm for Structured Bandits: A Sharp Characterization of Asymptotic Success / Failure*

### v1

@@ Abstract
We study the greedy (exploitation-only) algorithm in bandit problems with a known reward structure. We allow arbitrary finite reward structures, while prior work focused on a few specific ones. We fully characterize when the greedy algorithm asymptotically succeeds or fails, in the sense of sublinear vs. linear regret as a function of time. Our characterization identifies a partial identifiability property of the problem instance as the necessary and sufficient condition for the asymptotic success. Notably, once this property holds, the problem becomes easy—any algorithm will succeed (in the same sense as above), provided it satisfies a mild non-degeneracy condition. We further extend our characterization to contextual bandits and interactive decision-making with arbitrary feedback, and demonstrate its broad applicability across various examples.
@@ 1 Introduction
While exploration is central to bandit research, it presents significant challenges in practice, especially when an algorithm interacts with human users. First, exploration can be wasteful and risky for the current user, imposing a burden that may be considered unfair—or even unethical—since its benefits primarily accrue to future users. Second, exploration adds complexity to algorithm design, [^1] Most visibly, in contextual bandits with an “exploitation oracle” (Langford and Zhang, 2007; Agarwal et al., 2014). and its adoption in large-scale applications requires substantial buy-in and engineering support compared to a system that only exploits (Agarwal et al., 2016, 2017). Third, exploration may be incompatible with users’ incentives when actions are controlled by the users. For example, an online platform cannot force users to try and review new products; instead, users naturally gravitate toward well-reviewed or familiar options (Kremer et al., 2014). [^2] While exploration can be made incentive-compatible in such settings, doing so introduces additional cost and complexity (Kremer et al. (2014) and follow-up work, see Slivkins (2023) for an overview).
@@
However, beyond such examples, the broader picture remains unclear. Much research on bandits is devoted to problems with a known reward structure, e.g., linearity, Lipschitzness or convexity, where observing some actions provides useful information about others. Formally, a reward structure restricts the possible reward functions: functions from arms to their mean rewards. Reward structures reduce the need for explicit exploration, making the bandit problem more tractable. For some of them, $\mathtt{Greedy}$ in fact succeeds. A simple example: two-armed bandits with expected rewards that sum up to a known value. The literature provides examples of both failures and successes for a few specific reward structures, e.g., successes for linear contextual bandits (see Related Work). Likewise, large-scale experiments yield mixed results—some settings confirm the need for exploration, but others indicate that $\mathtt{Greedy}$ performs well even without it (Bietti et al., 2021). This contrast raises a fundamental question: When and why does $\mathtt{Greedy}$ succeed or fail?
Our Contributions. We study $\mathtt{Greedy}$ in bandits with general, known reward structures, finite but otherwise arbitrary. This formulation subsumes many structured bandit settings studied in prior work (possibly via discretization, see Section 2). We provide a complete characterization of when $\mathtt{Greedy}$ asymptotically succeeds (achieves sublinear regret) and when it fails (incurs linear regret). Our key insight is the identification of a new “partial identifiability” property of the problem instance, called self-identifiability, as a necessary and sufficient condition for the asymptotic success.
@@
Self-identifiability asserts that, given the reward structure, fixing the expected reward of a suboptimal arm uniquely identifies it as suboptimal. We prove that $\mathtt{Greedy}$ achieves sublinear regret under self-identifiability, and suffers from linear regret otherwise. The negative result is driven by the existence of a a decoy: informally, an alternative reward model such that its optimal arm is suboptimal for the true model and both models coincide on this arm. We show that with some positive probability, $\mathtt{Greedy}$ gets permanently stuck on such decoy, for an infinite time horizon.
@@
A crucial implication of our characterization is that the success of $\mathtt{Greedy}$ is not due to any particularly sophisticated behavior on its part but rather because self-identifiability makes the problem intrinsically “easy”. Indeed, we prove that whenever this condition holds, any reasonable algorithm (satisfying a mild non-degeneracy condition) also achieves sublinear regret.
We extend our characterization in two key directions. First, we generalize self-identifiability and the decoy construction to contextual bandits, showing that the same necessary and sufficient condition governs the asymptotic behavior of $\mathtt{Greedy}$ in this broader setting. Second, we consider a general framework of interactive decision-making with arbitrary auxiliary feedback, introduced in Foster et al. (2021), which includes bandit variants such as combinatorial semi-bandits as well as episodic reinforcement learning (see Section 5 for citations). We define an appropriate notion of $\mathtt{Greedy}$ for this setting, and derive a similar characterization.
@@
We illustrate our characterization with several examples, both positive and negative. We find that $\mathtt{Greedy}$ fails in linear bandits, Lipschitz bandits and “polynomial bandits” (with arms in $\mathbb{R}$ and polynomial expected rewards), and does so for almost all problem instances. For linear contextual bandits, $\mathtt{Greedy}$ succeeds if the context set is “sufficiently diverse”, but may fail if it is “low-dimensional”. For Lipschitz contextual bandits, $\mathtt{Greedy}$ behaves very differently, failing for almost all instances. One informal interpretation is that the success of $\mathtt{Greedy}$ requires both context diversity and a parametric reward structure.
@@
Discussion. We focus on the distinction between linear and sublinear regret, a standard “first-order” notion of success/failure for a bandit algorithm. Our positive results attain logarithmic worst-case regret rates, but possibly with a very large multiplicative constant determined by the reward structure. Our negative results establish a positive-constant (but possibly very small) probability of a “failure event” when $\mathtt{Greedy}$ gets permanently stuck on a decoy, for an infinite time horizon. Optimizing these “constants” for an arbitrary reward structure appears difficult. We make some progress in this direction, deriving an improved negative result for a special case.
@@
The greedy algorithm is initialized with some warm-up data collected from the same problem instance (and it needs at least 1 round of warm-up data to be well-defined). Our negative results require exactly one warm-up sample for each arm (for bandits) and one sample for each context-arm pair (for contextual bandits). All our positive results allow for an arbitrary amount of initial data. Thus, our characterization effectively defines “success” as sublinear regret for any amount of warm-up data, and “failure” as linear regret for some amount of warm-up data.
@@
Following the work on bandits with regression oracles (referenced below), we assume that $\mathtt{Greedy}$ is given a regression oracle: a subroutine to perform (least-squares) regression for the respective reward structure. Hence, we separate out computational issues, leveraging prior work on regression when and if it exists, and focus on the statistical guarantees.
@@
Related Work. Bandit reward structures studied in prior work include linear and combinatorial structures (e.g.,  Awerbuch and Kleinberg, 2008; McMahan and Blum, 2004; György et al., 2007; Cesa-Bianchi and Lugosi, 2012), convexity (e.g.,  Kleinberg, 2004; Flaxman et al., 2005; Bubeck et al., 2017), and Lipschitzness (e.g.,  Kleinberg, 2004; Kleinberg et al., 2008; Bubeck et al., 2011), as well as some others. Each of these is a long line of work on its own, with extensions to contextual bandits (e.g.,  Li et al., 2010; Slivkins, 2014). There’s also some work on bandits with arbitrary reward structures (Amin et al., 2011; Combes et al., 2017; Jun and Zhang, 2020; Degenne et al., 2020; Parys and Golrezaei, 2024), and particularly contextual bandits with regression oracles (e.g.,  Agarwal et al., 2012; Foster et al., 2018; Foster and Rakhlin, 2020; Simchi-Levi and Xu, 2022). Background on all this work can be found in books Slivkins (2019); Lattimore and Szepesvári (2020); Foster and Rakhlin (2023).
@@
For $\mathtt{Greedy}$, positive results with near-optimal regret rates focus on linear contextual bandits with diverse/smoothed contexts (Kannan et al., 2018; Bastani et al., 2021; Raghavan et al., 2023). The greedy algorithm is also known to attain $o(T)$ regret in various scenarios with a very large number of near-optimal arms (Bayati et al., 2020; Jedor et al., 2021). [^3] E.g., for Bayesian bandits with $\gg\sqrt{T}$ arms, where the arms’ mean rewards are sampled uniformly.
@@
Negative results for $\mathtt{Greedy}$ are derived for “non-structured” $K$-armed bandits: from trivial extensions of the single-sample-per-arm example mentioned above, to an exponentially stronger characterization of failure probability (Banihashem et al., 2023a), to various “near-greedy” algorithms / behaviors, both “frequentist” and “Bayesian” (same paper). Negative results for non-trivial reward structures concern dynamic pricing with linear demands (Harrison et al., 2012; den Boer and Zwart, 2014) and dynamic control in a generalized linear model (Lai and Robbins, 1982; Keskin and Zeevi, 2018). Banihashem et al. (2023a, b) also obtain negative results for the Bayesian version of $\mathtt{Greedy}$ in Bayesian bandits, under a certain “full support” assumption on the prior. [^4] Essentially, the prior allows all reward functions $\{\text{arms}\}\to[0,1]$, each with probability density at least $p>0$.
Map of the Paper. Section 3 focuses on multi-armed bandits, presenting our characterization in the most lucid way. Section 4 extends this characterization to contextual bandits. Section 5 is on $\mathtt{Greedy}$ for interactive decision-making with arbitrary feedback (along with all necessary setup). Examples are discussed in Section 6. Section 7 explains how self-identifiability makes the problem easy. While the body of the paper contains proof sketches, the detailed proof are in the appendix.
2 Preliminaries
@@
Let’s define the contextual bandit problem with a known reward structure, call it $\mathtt{StructuredCB}$. We have action set $\mathcal{A}$ and context set $\mathcal{X}$. In each round $t=1,2,\,\ldots$, a context $x_{t}\in\mathcal{X}$ arrives, an algorithm chooses an action (arm) $a_{t}\in\mathcal{A}$, and a reward $r_{t}\in\mathbb{R}$ is realized. The context is drawn independently from some fixed and known distribution over $\mathcal{X}$.  [^5] Whether the context distribution is known to the algorithm is inconsequential, since the greedy algorithm (particularly, the regression in Eq. (2.1)) does not use this knowledge. W.l.o.g., $\mathcal{X}$ is the support set of the context distribution. The reward $r_{t}$ is an independent draw from a unit-variance Gaussian with unknown mean $f^{*}(x_{t},a_{t})\in[0,1]$.  [^6] Gaussian reward noise is a standard assumption in bandits (along with e.g., 0-1 rewards), which we make for ease of presentation. Our positive results carry over to rewards with an arbitrary sub-Gaussian noise, without any modifications. Likewise, our negative results carry over to rewards $r_{t}\in[0,1]$ with an arbitrary near-uniform distribution, i.e., any distribution specified by a p.d.f. on $[0,1]$ such that the p.d.f. is bounded away from 00 by an absolute constant. A reward function is a function $f:\mathcal{X}\times\mathcal{A}\to[0,1]$; in particular, $f^{*}$ is the true reward function. The reward structure is given by a known class $\mathcal{F}$ of reward functions which contains $f^{*}$; the assumption $f^{*}\in\mathcal{F}$ is known as realizability. To recap, the problem instance is a pair $(f^{*},\mathcal{F})$, where $\mathcal{F}$ is known and $f^{*}$ is not.
@@
We focus on finite reward structures, i.e., assume (unless specified otherwise) that $\mathcal{X},\mathcal{A},\mathcal{F}$ are all finite. While this does not hold for most reward structures from prior work, one can discretize these reward structures to ensure finiteness. Indeed, when reward functions can take infinitely many values, one could consider some finite subset $S\subset[0,1]$, e.g., all integer multiples of some $\varepsilon>0$, and round each function value to the closest point in $S$. Likewise, one could discretize contexts, arms, or function parameters if needed, when they are represented as points in some metric space. [^7] A typical representation is real-valued vectors. One could take some finite point set $S$ in the metric space, and map each context (resp., arm or function parameter) to the closest point in $S$. Alternatively, one could define finite reward structures directly, with similar discretizations built into their respective definitions (see Section 6 for several examples).
@@
Regret at round $t$ is defined in a standard way: $R(t):=\sum_{s\in[t]}\left(\,r^{*}(x_{s})-r_{s}\,\right)$, where $r^{*}(x):=\max_{a\in\mathcal{A}}f^{*}(x,a)$ is the best expected reward for a given context $x$. We are interested in expected regret $\operatornamewithlimits{\mathbb{E}}\left[\,R(t)\,\right]$ as a function of time $t$.
The greedy algorithm. The greedy bandit algorithm ($\mathtt{Greedy}$), is defined as follows.
@@
It is initialized with $T_{0}\geq 1$ rounds of warm-up data, denoted $t\in[T_{0}]$.  [^8] We also refer to the first $T_{0}$ rounds as warm-up stage, and the subsequent rounds as main stage. Each such round yields a context-arm pair $(x_{t},a_{t})\in\mathcal{X}\times\mathcal{A}$ chosen exogenously, and reward $r_{t}\in\mathbb{R}$ drawn independently from the resp. reward distribution: unit-variance Gaussian with mean $f^{*}(x_{t},a_{t})$.
At each round $t>T_{0}$, $\mathtt{Greedy}$ computes a reward function via least-squares regression:
@@
We posit a “regression oracle” which implements the regression, as discussed in Section 1. Note that there are no ties in (2.1) with probability one over the random rewards. Once reward function $f_{t}$ is chosen, the algorithm chooses the best arm for $f_{t}$ and context $x_{t}$, i.e.,
@@
Structured bandits. Multi-armed bandits with a known structure is a paradigmatic special case when there is only one context, $|\mathcal{X}|=1$; call it $\mathtt{StructuredMAB}$. Then the context can be suppressed from the notation; e.g., reward functions map arms to $[0,1]$. An arm is called optimal for a given reward function $f$ (or, by default, for $f=f^{*}$) if it maximizes expected reward $f(\cdot)$, and suboptimal otherwise.
@@
3 Characterization for multi-armed bandits
We focus on $\mathtt{StructuredMAB}$: $K$-armed bandits with a known reward structure $\mathcal{F}$. We start with two key definitions. Self-identifiability (which drives the positive result) asserts that fixing the expected reward of any suboptimal arm identifies this arm as suboptimal.

### v2

@@ Abstract
We study the greedy (exploitation-only) algorithm in bandit problems with a known reward structure. We allow arbitrary finite reward structures, while prior work focused on a few specific ones. We fully characterize when the greedy algorithm asymptotically succeeds or fails, in the sense of sublinear vs. linear regret as a function of time. Our characterization identifies a partial identifiability property of the problem instance as the necessary and sufficient condition for the asymptotic success. Notably, once this property holds, the problem becomes easy—any algorithm will succeed (in the same sense as above), provided it satisfies a mild non-degeneracy condition. Our characterization extends to contextual bandits and interactive decision-making with arbitrary feedback. Examples demonstrating broad applicability and extensions to infinite reward structures are provided.
@@
Version history. First version: March 2025. This version: May 2025 (adds Section 6 on infinite function classes).
@@ 1 Introduction
While exploration is central to bandit research, it presents significant challenges in practice, esp. when an algorithm interacts with human users. First, exploration can be wasteful and risky for the current user, imposing a burden that may be considered unfair since its benefits primarily accrue to future users. Second, exploration adds complexity to algorithm design,and its adoption in large-scale applications requires substantial buy-in and engineering support compared to a system that only exploits (Agarwal et al., 2016, 2017). Third, exploration may be incompatible with users’ incentives when actions are controlled by the users. E.g., an online platform cannot force users to try and review new products; instead, users gravitate toward well-reviewed or familiar options (Kremer et al., 2014). [^1] While exploration can be made incentive-compatible in such settings, doing so introduces additional cost and complexity (Kremer et al. (2014) and follow-up work, see Slivkins (2023) for an overview).
@@
However, beyond such examples, the broader picture remains murky, especially for the widely-studied structured bandits—bandit problems with a known reward structure (e.g., linearity, Lipschitzness, convexity)—where observing some actions provides useful information about others. Formally, a reward structure restricts the possible reward functions that map arms to their mean rewards. Reward structures reduce the need for explicit exploration, making the bandit problem more tractable. For some of them, $\mathtt{Greedy}$ in fact succeeds, e.g., two-armed bandits with expected rewards that sum up to a known value. The literature provides a few examples of failure for some specific (one-dimensional, linear) reward structures, and a few non-trivial examples of success (e.g., for linear contextual bandits), more on this in Related Work. Likewise, large-scale experiments yield mixed results: some settings confirm the need for exploration, but others indicate that $\mathtt{Greedy}$ performs well even (Bietti et al., 2021). This contrast raises a fundamental question: When—and why—does $\mathtt{Greedy}$ fail or succeed?
Our Contributions. We work towards the missing foundation for structured bandits: a general theory of $\mathtt{Greedy}$. Our main result allows finite, but otherwise arbitrary reward structures. We provide a complete characterization of when $\mathtt{Greedy}$ asymptotically fails (incurs linear regret) vs when it succeeds (achieves sublinear regret). Our characterization applies to every problem instance, resolving it in the positive or negative direction, not (just) in the worst case over a particular reward structure. The negative results are of primary interest here, as they substantiate the common belief that $\mathtt{Greedy}$ performs poorly, and the positive results serve to make the characterization precise.
@@
A key insight is identifying a new “partial identifiability” property of the problem instance, called self-identifiability, as a necessary and sufficient condition for the asymptotic success. Self-identifiability asserts that, given the reward structure, fixing the expected reward of a suboptimal arm uniquely identifies it as suboptimal. We prove that $\mathtt{Greedy}$ achieves sublinear regret under self-identifiability, and suffers from linear regret otherwise. The negative result is driven by the existence of a decoy: informally, an alternative reward model such that its optimal arm is suboptimal for the true model and both models coincide on this arm. We show that with some positive probability, $\mathtt{Greedy}$ gets permanently stuck on such decoy, for an infinite time horizon.
@@
$\mathtt{Greedy}$ succeeds (only) because self-identifiability makes the problem instance intrinsically “easy”: any reasonable algorithm (under a mild non-degeneracy condition) achieves sublinear regret.
Our characterization allows for essentially an arbitrary interaction protocol between the algorithm and the environment (Section 5). Specifically, we handle interactive decision-making with arbitrary auxiliary feedback (Foster et al., 2021), which subsumes contextual bandits and combinatorial semi-bandits, as well as episodic reinforcement learning. That said, before moving to this full generality, our presentation focuses on contextual bandits, where we obtain quantitatively stronger guarantees (Section 4), and “vanilla” bandits as a paradigmatic case for building key intuition (Section 3).
The second main result of this work is a partial characterization which handles arbitrary infinite (e.g., continuous) reward structures. It applies to structured bandits with finite action sets, requiring stronger notions of self-identifiability and decoy existence, as well as new analysis ideas, see Section 6.
@@
We apply our machinery to several examples, both positive and negative (Section 7). We find that $\mathtt{Greedy}$ fails in linear bandits, Lipschitz bandits and “polynomial bandits” (with arms in $\mathbb{R}$ and polynomial expected rewards), and does so for almost all problem instances. For linear contextual bandits, $\mathtt{Greedy}$ succeeds if the context set is “sufficiently diverse”, but may fail if it is “low-dimensional”. For Lipschitz contextual bandits, $\mathtt{Greedy}$ behaves very differently, failing for almost all instances. One informal takeaway is that $\mathtt{Greedy}$ fails as a common case for most/all bandit structures of interest, whereas for contextual bandits it can go either way. The success of $\mathtt{Greedy}$ appears to require context diversity and a parametric reward structure.
Discussion. While our first main result—the complete characterization—assumes finite reward structures, most infinite structures of interest admit meaningful finite analogs via discretization (see Section 2 and illustrations in Section 7). Moreover, we relax the finiteness assumption in Section 6.
@@
The linear vs sublinear regret is a standard “first-order” notion of success/failure in bandits. Our positive results attain logarithmic worst-case regret rates, possibly with a large multiplicative constant determined by the reward structure. Our negative results establish a positive-constant (but possibly very small) probability of a “failure event” when $\mathtt{Greedy}$ gets permanently stuck on a decoy, for an infinite time horizon. Optimizing these constants for an arbitrary reward structure appears difficult. However, we achieve much better constants for the partial characterization in Section 6.
@@
The greedy algorithm is initialized with some warm-up data collected from the same problem instance (and it needs at least $1$ warm-up sample to be well-defined). Our negative results require exactly one warm-up sample for each context-arm pair. All our positive results allow for an arbitrary amount of initial data. Thus, our characterization effectively defines “success” as sublinear regret for any amount of warm-up data, and “failure” as linear regret for some amount of warm-up data.
@@
We assume that $\mathtt{Greedy}$ is given a regression oracle: a subroutine to perform (least-squares) regression given the reward structure. As in “bandits with regression oracles” (referenced below), we separate out computational issues, leveraging prior work on regression, and focus on the statistical guarantees.
@@
Related Work. Bandit reward structures studied in prior work include linear and combinatorial structures (e.g.,  Awerbuch and Kleinberg, 2008; McMahan and Blum, 2004; György et al., 2007; Cesa-Bianchi and Lugosi, 2012), convexity (e.g.,  Kleinberg, 2004; Flaxman et al., 2005; Bubeck et al., 2017), and Lipschitzness (e.g.,  Kleinberg, 2004; Kleinberg et al., 2008; Bubeck et al., 2011), as well as some others. Each of these is a long line of work on its own, with extensions to contextual bandits (e.g.,  Li et al., 2010; Slivkins, 2014). There’s also some work on bandits with arbitrary reward structures (Amin et al., 2011; Combes et al., 2017; Jun and Zhang, 2020; Degenne et al., 2020; Parys and Golrezaei, 2024), and particularly contextual bandits with regression oracles (e.g.,  Agarwal et al., 2012; Foster et al., 2018; Foster and Rakhlin, 2020; Simchi-Levi and Xu, 2022). For more background, see books Slivkins (2019); Lattimore and Szepesvári (2020); Foster and Rakhlin (2023).
@@
For $\mathtt{Greedy}$, positive results with near-optimal regret rates focus on linear contextual bandits with diverse/smoothed contexts (Kannan et al., 2018; Bastani et al., 2021; Raghavan et al., 2023; Kim and Oh, 2024). The greedy algorithm is also known to attain $o(T)$ regret in various scenarios with a very large number of near-optimal arms (Bayati et al., 2020; Jedor et al., 2021). [^2] E.g., for Bayesian bandits with $\gg\sqrt{T}$ arms, where the arms’ mean rewards are sampled uniformly.
@@
Negative results for $\mathtt{Greedy}$ are derived for “non-structured” $K$-armed bandits: from trivial extensions of the single-sample-per-arm example mentioned above, to an exponentially stronger characterization of failure probability (Banihashem et al., 2023a), to various “near-greedy” algorithms / behaviors, both “frequentist” and “Bayesian” (same paper). Negative results for non-trivial reward structures concern dynamic pricing with linear demands (Harrison et al., 2012; den Boer and Zwart, 2014) and dynamic control in a generalized linear model (Lai and Robbins, 1982; Keskin and Zeevi, 2018). Banihashem et al. (2023a, b) also obtain negative results for the Bayesian version of $\mathtt{Greedy}$ in Bayesian bandits, under a certain “full support” assumption on the prior. [^3] Essentially, the prior allows all reward functions $\{\text{arms}\}\to[0,1]$ with probability density at least $p>0$.
2 Preliminaries: structured contextual bandits ($\mathtt{StructuredCB}$)
@@
We have action set $\mathcal{A}$ and context set $\mathcal{X}$. In each round $t=1,2,\,\ldots$, a context $x_{t}\in\mathcal{X}$ arrives, an algorithm chooses an action (arm) $a_{t}\in\mathcal{A}$, and a reward $r_{t}\in\mathbb{R}$ is realized. The context is drawn independently from some fixed and known distribution over $\mathcal{X}$.  [^4] Whether the context distribution is known to the algorithm is inconsequential, since $\mathtt{Greedy}$ (particularly, the regression in Eq. (2.1)) does not use this knowledge. W.l.o.g., $\mathcal{X}$ is the support set of the context distribution. The reward $r_{t}$ is an independent draw from a unit-variance Gaussian with unknown mean $f^{*}(x_{t},a_{t})\in[0,1]$.  [^5] Gaussian reward noise is a standard assumption in bandits (along with e.g., 0-1 rewards), which we make for ease of presentation. Our positive results carry over to rewards with an arbitrary sub-Gaussian noise, without any modifications. Likewise, our negative results carry over to rewards $r_{t}\in[0,1]$ with an arbitrary near-uniform distribution, i.e., one specified by a p.d.f. on $[0,1]$ which is bounded away from 00 by an absolute constant. A reward function is a function $f:\mathcal{X}\times\mathcal{A}\to[0,1]$; in particular, $f^{*}$ is the true reward function. The reward structure is given by a known class $\mathcal{F}$ of reward functions which contains $f^{*}$; the assumption $f^{*}\in\mathcal{F}$ is known as realizability. To recap, the problem instance is a pair $(f^{*},\mathcal{F})$, where $\mathcal{F}$ is known and $f^{*}$ is not.
@@
We focus on finite reward structures, i.e., assume (unless specified otherwise) that $\mathcal{X},\mathcal{A},\mathcal{F}$ are all finite. While this does not hold for most reward structures from prior work, one can discretize them to ensure finiteness. Indeed, when reward functions can take infinitely many values, one could round each function value to the closest point in some finite subset $S\subset[0,1]$, e.g., all integer multiples of some $\varepsilon>0$. Likewise, one could discretize contexts, arms, or function parameters, when they are represented as points in some metric space, e.g., as real-valued vectors. Or, one could define finite reward structures directly, with similar discretizations built-in (see Section 7 for examples).
@@
We are interested in expected regret $\operatornamewithlimits{\mathbb{E}}\left[\,R(t)\,\right]$ as a function of round $t$. Regret is standard: $R(t):=\sum_{s\in[t]}\left(\,r^{*}(x_{s})-r_{s}\,\right)$, where $r^{*}(x):=\max_{a\in\mathcal{A}}f^{*}(x,a)$, best expected reward given context $x$.
@@
The greedy algorithm ($\mathtt{Greedy}$) is defined as follows. It is initialized with $T_{0}\geq 1$ rounds of warm-up data, denoted $t\in[T_{0}]$.  [^6] We also refer to the first $T_{0}$ rounds as warm-up stage, and the subsequent rounds as main stage. Each such round yields a context-arm pair $(x_{t},a_{t})\in\mathcal{X}\times\mathcal{A}$ chosen exogenously, and reward $r_{t}\in\mathbb{R}$ drawn independently from the resp. reward distribution: unit-variance Gaussian with mean $f^{*}(x_{t},a_{t})$. At each round $t>T_{0}$, $\mathtt{Greedy}$ computes a reward function via least-squares regression (implemented via a “regression oracle”, as per Section 1):
@@
Note that there are no ties in (2.1) with probability one over the random rewards. Once reward function $f_{t}$ is chosen, the algorithm chooses the best arm for $f_{t}$ and context $x_{t}$, i.e.,
@@
3 Characterization for structured bandits
Let us focus on the paradigmatic special case of multi-armed bandits, call it $\mathtt{StructuredMAB}$. Formally, there is only one context, $|\mathcal{X}|=1$. The context can be suppressed from the notation; e.g., reward functions map arms to $[0,1]$. An arm is called optimal for a given reward function $f$ (or, by default, for $f=f^{*}$) if it maximizes expected reward $f(\cdot)$, and suboptimal otherwise.
We start with two key definitions. Self-identifiability (which drives the positive result) asserts that fixing the expected reward of any suboptimal arm identifies this arm as suboptimal.